# tableReferences for Table S1

### v1

@@ Abstract
Generative models are invaluable in many fields of science because of their ability to capture high-dimensional and complicated distributions, such as photo-realistic images, protein structures, and connectomes. How do we evaluate the samples these models generate? This work aims to provide an accessible entry point to understanding popular notions of statistical distances, requiring only foundational knowledge in mathematics and statistics. We focus on four commonly used notions of statistical distances representing different methodologies: Using low-dimensional projections (Sliced-Wasserstein; SW), obtaining a distance using classifiers (Classifier Two-Sample Tests; C2ST), using embeddings through kernels (Maximum Mean Discrepancy; MMD), or neural networks (Fréchet Inception Distance; FID). We highlight the intuition behind each distance and explain their merits, scalability, complexity, and pitfalls. To demonstrate how these distances are used in practice, we evaluate generative models from different scientific domains, namely a model of decision making and a model generating medical images. We showcase that distinct distances can give different results on similar data. Through this guide, we aim to help researchers to use, interpret, and evaluate statistical distances for generative models in science.
@@ 1 Introduction
Generative models that produce samples of complex, high-dimensional data, have recently come to the forefront of public awareness due to their utility in a variety of scientific, clinical, engineering, and commercial domains (Bond-Taylor et al., 2021). Prominent examples include StableDiffusion and DALL-E for generating photo-realistic images (Rombach et al., 2022), WaveNet (Oord et al., 2016) for audio synthesis, and Generative Pre-trained Transformer (GPT;  Radford et al. 2018; 2019, Brown et al. 2020) for text generation. Besides this new wave of generative models, different scientific disciplines have a long history of building data generating models which capture specific processes. For example in neuroscience, the occurrence of action potentials is modeled at all different levels of detail (e.g. single neuron voltage dynamics (Hodgkin & Huxley, 1952) or at a phenomenological level (Pillow et al., 2008)), whereas in e.g. astrophysics there exist various models to simulate galaxy formation (Somerville & Davé, 2015). Along generating novel synthetic samples, generative models can be leveraged for specific tasks, such as sample generation conditioned on class labels (e.g., diseased vs. healthy brain scans, molecules that can or cannot be synthesized; Urbina et al. 2022, class-conditional image generation; van den Oord et al. 2016, Dockhorn et al. 2022, forecasting future states of a dynamical system; Durstewitz et al. 2023, Jacobs et al. 2023, Brenner et al. 2022), data imputation (e.g., Vetter et al. 2023, Lugmayr et al. 2022), data augmentation for downstream tasks (Rommel et al., 2022), and many more (see also Table S1).
@@
These powerful capabilities are enabled by the premise that generative models accurately learn a high-dimensional distribution from which we assume our dataset was sampled. The dimensions can correspond to anything from individual pixels or graphs to arbitrary features of physical or abstract objects. When aiming to build generative models that better capture the true underlying data distribution, we need to answer a key question: How accurately does our generative model mimic the true data distribution?
@@
Manual inspection of generated samples can be a good first check, e.g., in image or audio generation, where we can directly assess the visual likeness or sound quality of the samples (Gerhard et al., 2013, Vallez et al., 2022, Jayasumana et al., 2023). In general, however, we would like to have quantitative distances to measure the similarity of distributions, for instance to compare and benchmark different models. Many measures have been proposed that provide a quantitative way to assess the similarity of two distributions based on various aspects of their moments, spread, central tendency, and overall probability density (Fig. 1). Some of these measures require likelihood evaluations, as is possible with generative models such as Gaussian Mixture Models, Normalizing Flows, Variational Autoencoders, autoregressive models or diffusion models (Bishop, 2006, Papamakarios et al., 2021, Box et al., 2015, Kingma & Welling, 2014, Yenduri et al., 2023, Ho et al., 2020, Song et al., 2021). However, many contemporary machine learning models (e.g., Generative Adversarial Networks and Energy-Based Models; Goodfellow et al. 2014, Rezende et al. 2014, Hinton et al. 2006) and scientific simulators (e.g., single neuron voltage dynamics; Hodgkin & Huxley 1952) only define the likelihood implicitly, i.e. we can not explicitly evaluate their likelihood. In this work we focus on distances that can be applied to both classes of models, i.e. distances that can be computed only based on generated samples.
@@
Figure 1: The need for statistical distances in scientific generative modeling. (a) An example target distribution, $p_{true}(x)$, and two learned distributions ($p_{1}(x)$ and $p_{2}(x)$) of different models trained to capture $p_{true}(x)$. All three distributions share the same mean and marginal variances, despite having distinct shapes. However, an appropriate sample-based distribution distance $D$ can determine that $p_{2}(x)$ is more similar to $p_{true}(x)$. (b) Scientific applications often require evaluating high-dimensional distributions, such as distributions for images or tabular data. In this example, each point represents an X-ray image, where each dimension is one pixel.
Here, we provide a guide to understanding commonly used sample-based statistical distances. Note that with distance, we do not necessarily refer to a distance metric in the mathematical sense (i.e., satisfying symmetry and the triangle inequality) but to a general measure of dissimilarity between two distributions. Our goal is not to provide a comprehensive review of statistical distances, as there are already a number of excellent resources for that purpose, especially in specific domains of application (Borji, 2019, Xu et al., 2018, Lopez-Paz & Oquab, 2016, Lueckmann et al., 2021). We refer readers to those works for a deeper dive into mathematical properties and empirical comparisons  (Cox et al., 1984, Gibbs & Su, 2002, Basseville, 2013, Cai & Lim, 2022, Muandet et al., 2017, Theis et al., 2016, Betzalel et al., 2022). In this guide, we instead focus on four commonly applied sample-based distances in the machine learning literature for evaluating eventually high-dimensional generative models. They represent different methodologies for defining statistical distance: Using low-dimensional projections (Sliced-Wasserstein; SW), obtaining a distance using classifiers (Classifier Two-Sample Tests; C2ST), using embeddings through kernels (Maximum Mean Discrepancy; MMD) or neural networks (Fréchet Inception Distance; FID). We aim to provide an intuition for how and when to apply these distances, and to build a solid foundation for navigating the extensive literature on statistical distances.
@@
Towards this goal, we provide for each of the four distances an intuitive and graphical explanation (Section 2). We then perform a systematic evaluation of their robustness as a function of dataset size, data dimensionality, and other factors, such as data multimodality (Section 3). Finally, in Section 4, we demonstrate how these distances can be applied to compare generative models in different scientific domains: We evaluate low dimensional models of decision making in behavioral neuroscience and generative models of medical X-ray images. We show the importance of using multiple complementary distances, as distinct distances can give different results when comparing the same sets of samples. By presenting these resources, we aim to empower researchers to choose, implement, and evaluate the use and outcomes of statistical distances for generative models in science.
@@ 2.1 Slicing-based: Sliced-Wasserstein (SW) distance
Computing distance between distributions suffers from the curse of dimensionality, where the computational cost of computing the distance increases very rapidly as the dimensionality of the data increases. This problem is especially restricting when the distance is used as part of a loss function in optimization problems, since in this case it needs to be evaluated many times. This has prompted the notion of “sliced” distances, which have become increasingly popular in recent years (Kolouri et al., 2019, Nadjahi et al., 2020, Goldfeld & Greenewald, 2021). The main idea behind slicing is that for many existing statistical distances we can efficiently evaluate the distance in low-dimensional spaces, especially in one dimension. Therefore, the “slices” are typically one-dimensional lines through the data space (Fig. 2a). All data points from each distribution are projected onto this line by finding their nearest point on the line, giving a one-dimensional distribution of projected data points (Fig. 2b). The distance measure of interest between the resulting one-dimensional distributions can then be computed efficiently. However, computing the random projection could lead to an unreliable measure of distance: Distinct distributions can produce the same one-dimensional projections. Therefore, we repeat the slicing process for many different slices and average the resulting distances. More formally, we compute the expected distance in one dimension between the projections of the respective distributions onto (uniformly) random directions on the unit sphere. As long as the distance of choice is a valid metric in one dimension, the sliced distance defined in this way is guaranteed to be a valid metric as well (Nadjahi et al., 2020, Proposition 1 (iii)). The most popular example of a sliced distance metric is the Sliced-Wasserstein (SW) distance (Fig. 2). However, we note that slicing has also been done for other distance measures, such as MMD with a specific choice of kernel (Hertrich et al., 2023) and mutual information (Goldfeld & Greenewald, 2021).
Figure 3: Computing Wasserstein distance. Two transport maps mapping the samples from a distribution $p_{1}$ (black) to samples from another distribution $p_{2}$ (blue), shown by arrows. The color of the arrow corresponds to the cost (Euclidean distance) between $x_{i}$ and $y_{i}$. (a) Randomly chosen transport map. (b) The optimal transport map, giving the smallest total cost. The total cost for the optimal map in (b) is the Wasserstein distance between these two sets of samples.
Wasserstein or earth mover distance
@@
Intuitively, if two given probability distributions are thought of as two piles of dirt, the Wasserstein distance measures the (minimal) cost of “transporting” one pile of dirt to another (Panaretos & Zemel, 2019). The formal definition of the Wasserstein distance for continuous distributions, derived from optimal transport, is described in Section A.3. Here, we provide a more intuitive definition given a fixed set of samples from two distributions. Suppose we have samples $\{x_{1},\ldots,x_{N}\}\subset\mathbb{R}^{d}$ sampled from a distribution $p_{1}$ and $\{y_{1},...,y_{N}\}\subset\mathbb{R}^{d}$ sampled from another distribution $p_{2}$. Given any distance metric between two vectors in $\mathbb{R}^{d}$, $D(\cdot,\cdot)$, we can construct the cost matrix $C$, as the matrix of pairwise distances between the samples $x_{i}$ and $y_{j}$:
@@
Recalling the earth-mover distance analogy, we want to map each $x_{i}$ to exactly one $y_{j}$, in such a way that the cost of doing so is minimized. The minimum transport map is then defining the Wasserstein distance (for the metric $D$) between the two empirical distributions. Throughout this work, we use the commonly used Euclidean metric, $L^{2}$, leading to the Wasserstein-2 and Sliced Wasserstein-2 distances. More precisely, we define a “transport map” to be a permutation matrix, $\pi\in\{0,1\}^{N\times N}$, which is a matrix with exactly one nonzero entry in each row. The entry $\pi_{ij}=1$ means that we transport the point $x_{i}$ to the point $y_{j}$. Then finding the transport map that minimizes the overall cost can be stated as
@@
Solving the optimal transport problem (Eq. (2)) with the Hungarian method has a time complexity of $O(N^{3})$ in the number of samples $N$ (although faster approximations exists, see Peyré et al. 2017). However, in the special case where the data is one-dimensional, the Wasserstein distance can be calculated by sorting the two datasets, obtaining the order statistics $\{x_{(1)},..,x_{(N)}\}$ and $\{y_{(1)},...,y_{(N)}\}$ and computing the sum of the distances $\sum_{i}D(x_{(i)},y_{(i)})$. This has a time complexity of $O(N\log(N))$. Thus, slicing the Wasserstein distance with one-dimensional projections becomes very efficient. While the value of the SW distance does not converge to the Wasserstein distance, even in the case of an infinite number of data samples and slices, the SW distance is a metric (in the mathematical sense) as long as $D$ is a metric on $\mathbb{R}^{d}$ and it acts as a lower bound to the Wasserstein distance (Nadjahi, 2021).
The Wasserstein distance and its sliced variant have several attractive properties: they can be computed differentiably; their computations do not rely heavily on choices of hyperparameters; and the sliced variant is very fast to compute. However, a disadvantage of the Wasserstein distance is its transparency: The numerical value of the Wasserstein distance has no intuitive interpretation due to its definition in terms of optimal transport maps. Additionally, per definition, the sliced variant is insensitive to differences within orthogonal subspaces of the slices. To still capture differences in all dimensions, naively, one would have to increase the number of slices (in the worst case) exponentially with the dimension, diminishing computational efficiency. However, other approaches exist to reduce this problem for nonlinear (Kolouri et al., 2019) or other specific (Deshpande et al., 2019; 2018) slices. Furthermore, slicing may also be relaxed to other kinds of data specific projections, such as Fourier features for stationary time series or locally connected projections for images (Du et al., 2023, Cazelles et al., 2020).

### v2

@@ Abstract
Generative models are invaluable in many fields of science because of their ability to capture high-dimensional and complicated distributions, such as photo-realistic images, protein structures, and connectomes. How do we evaluate the samples these models generate? This work aims to provide an accessible entry point to understanding popular sample-based statistical distances, requiring only foundational knowledge in mathematics and statistics. We focus on four commonly used notions of statistical distances representing different methodologies: Using low-dimensional projections (Sliced-Wasserstein; SW), obtaining a distance using classifiers (Classifier Two-Sample Tests; C2ST), using embeddings through kernels (Maximum Mean Discrepancy; MMD), or neural networks (Fréchet Inception Distance; FID). We highlight the intuition behind each distance and explain their merits, scalability, complexity, and pitfalls. To demonstrate how these distances are used in practice, we evaluate generative models from different scientific domains, namely a model of decision-making and a model generating medical images. We showcase that distinct distances can give different results on similar data. Through this guide, we aim to help researchers to use, interpret, and evaluate statistical distances for generative models in science.
@@ 1 Introduction
Generative models that produce samples of complex, high-dimensional data, have recently come to the forefront of public awareness due to their utility in a variety of scientific, clinical, engineering, and commercial domains (Bond-Taylor et al., 2021). Prominent examples include StableDiffusion (SD) and DALL-E for generating photo-realistic images (Rombach et al., 2022a), WaveNet (Oord et al., 2016) for audio synthesis, and Generative Pre-trained Transformer (GPT; Radford et al. 2018; 2019, Brown et al. 2020) for text generation. In addition to the recent surge in generative models, many scientific disciplines have a long history of developing data-generating models that capture specific processes. In neuroscience, for example, the occurrence of action potentials is commonly modeled at varying levels of detail (e.g., single neuron voltage dynamics; Hodgkin & Huxley 1952, or at a phenomenological level; Pillow et al. 2008), whereas in, e.g., astrophysics there exist various models to simulate galaxy formation (Somerville & Davé, 2015). Along with generating novel synthetic samples, generative models can be leveraged for specific tasks, such as sample generation conditioned on class labels (e.g., diseased vs. healthy brain scans Pombo et al. 2021, molecules that can or cannot be synthesized; Urbina et al. 2022, class-conditional image generation; van den Oord et al. 2016, Dockhorn et al. 2022), forecasting future states of a dynamical system (e.g.,  Jacobs et al. 2023, Brenner et al. 2022, Pals et al. 2024), generating neural population activity conditioned on visual stimuli or behavior (e.g., Molano-Mazon et al. 2018, Bashiri et al. 2021, Schulz et al. 2024, Kapoor et al. 2024), data imputation (e.g., Vetter et al. 2023, Lugmayr et al. 2022), data augmentation for downstream tasks (Rommel et al., 2022), and many more (see also Table S1).
@@
These powerful capabilities are enabled by the premise that generative models can produce samples from the high-dimensional distribution from which we assume our dataset was sampled. The dimensions can correspond to anything from individual pixels or graphs to arbitrary features of physical or abstract objects. When aiming to build generative models that better capture the true underlying data distribution, we need to answer a key question: How accurately do samples from our generative model mimic those from the true data distribution?
@@
Manual inspection of generated samples can be a good first check, e.g., in image or audio generation, where we can directly assess the visual likeness or sound quality of the samples (Gerhard et al., 2013, Vallez et al., 2022, Jayasumana et al., 2023). In general, however, we would like to quantitatively compare the similarity of distributions, for instance to benchmark different generative models. Many measures have been proposed that assess the similarity of two distributions based on various aspects of their moments or the overall probability density (Fig. 1). Some of these measures require likelihood evaluations, as is possible with generative models such as Gaussian Mixture Models, Normalizing Flows, Variational Autoencoders, autoregressive models or diffusion models (Bishop, 2006, Papamakarios et al., 2021, Box et al., 2015, Kingma & Welling, 2014, Yenduri et al., 2024, Ho et al., 2020, Song et al., 2021). However, many contemporary machine learning models (e.g., Generative Adversarial Networks and Energy-Based Models; Goodfellow et al. 2014, Rezende et al. 2014, Hinton et al. 2006) and scientific simulators (e.g., single neuron voltage dynamics; Hodgkin & Huxley 1952) only define the likelihood implicitly, i.e., we can not explicitly evaluate their likelihood. Statistical distances that can be computed based on samples only are therefore invaluable for comparing both classes of generative models (and to real data) in scientific contexts.
@@
Figure 1: The need for statistical distances in scientific generative modeling. (a) An example target distribution, $p_{true}(x)$, and two learned distributions ($p_{1}(x)$ and $p_{2}(x)$) of different models trained to capture $p_{true}(x)$. All three distributions share the same mean and marginal variances, despite having distinct shapes. However, an appropriate sample-based distribution distance $D$ can determine that $p_{2}(x)$ is more similar to $p_{true}(x)$. (b) Scientific applications often require evaluating high-dimensional distributions, such as distributions of images or tabular data. In this example, each point represents an X-ray image, where each dimension is one pixel.
Here, we present a guide that aims to serve as an accessible entry point to understanding commonly used sample-based statistical distances. Towards this goal, we provide explanations, comparisons, and example applications of four commonly applied distances: Sliced-Wasserstein (SW), Classifier Two-Sample Tests (C2ST), Maximum Mean Discrepancy (MMD), and Fréchet Inception Distance (FID). With these resources, we aim to empower researchers to choose, implement, and evaluate the usage and outcomes of statistical distances for generative models in science. Note that, with distance, we do not necessarily refer to a distance metric in the mathematical sense (i.e., satisfying symmetry and the triangle inequality) but to a general measure of dissimilarity between two distributions (however, some considered distances are in fact metrics).
The general and didactic nature of this guide means it can neither be comprehensive nor provide a clear-cut answer as to which distance is ‘best’, as there are numerous choices, and the ‘ideal’ one strongly depends on the domain of application. Previous articles have provided useful and extensive reviews for specific use cases. For example, Theis et al. (2016) discusses commonly used criteria for evaluating generative models of images, Borji (2019), Xu et al. (2018), Yang et al. (2023) compare metrics specific to evaluating GANs (including, e.g., specialized variants of the FID), Basseville (2013) provides an extensive overview of previous works on divergences, and Gibbs & Su (2002) analyze the theoretical relationships among ‘classic’ distances (including, e.g., Wasserstein and KL-divergence). Rather, by going through four different classes of sample-based distances in detail and systematically comparing them on synthetic and real-world applications, we aim to provide a solid foundation for navigating the extensive literature on statistical distances, and to enable readers to reason about other related distances not covered here.
@@
The outline of this guide is as follows: First, we provide an intuitive and graphical explanation for each of the four distances (Section 2). We then perform a systematic evaluation of their robustness as a function of dataset size, data dimensionality, and other factors, such as data multimodality (Section 3). Finally, we demonstrate how these distances can be applied to compare generative models in different scientific domains (Section 4): We evaluate low dimensional models of decision making in behavioral neuroscience and generative models of medical X-ray images. We show the importance of using multiple complementary distances, as distinct distances can give different results when comparing the same sets of samples.
@@ 2.1 Slicing-based: Sliced-Wasserstein (SW) distance
Computing distance between distributions suffers from the curse of dimensionality, where the computational cost of computing the distance increases very rapidly as the dimensionality of the data increases. This problem is especially restricting when the distance is used as part of a loss function in optimization problems, since in this case it must be evaluated many times. This has prompted the notion of “sliced” distances, which have become increasingly popular in recent years (Kolouri et al., 2019, Nadjahi et al., 2020, Goldfeld & Greenewald, 2021). The main idea behind slicing is that for many existing statistical distances we can efficiently evaluate the distance in low-dimensional spaces, especially in a single dimension. Therefore, the “slices” are typically one-dimensional lines through the data space (Fig. 2a). All data points from each distribution are projected onto this line by finding their nearest point on the line, giving a one-dimensional distribution of projected data points (Fig. 2b). The distance measure of interest between the resulting one-dimensional distributions can then be computed efficiently. However, computing the random projection could lead to an unreliable measure of distance as distinct distributions can produce the same one-dimensional projections. Therefore, we repeat the slicing process for many different slices and average the resulting distances. More formally, we compute the expected distance in one dimension between the projections of the respective distributions onto (uniformly) random directions on the unit sphere. As long as the distance of choice is a valid metric in one dimension, the sliced distance defined in this way is guaranteed to be a valid metric as well (Nadjahi et al., 2020, Proposition 1 (iii)). The most popular example of a sliced distance metric is the Sliced-Wasserstein (SW) distance (Fig. 2). The Wasserstein distance and its sliced variant have several attractive properties: They can be computed differentiably; their computations do not rely heavily on choices of hyperparameters; and the sliced variant is very fast to compute. However, we note that slicing has also been done for other distance measures, such as MMD with a specific choice of kernel (Hertrich et al., 2024) and mutual information (Goldfeld & Greenewald, 2021). We provide a formal definition of the Wasserstein distance below, and of the Sliced-Wasserstein distance in Appendix A.3.
Definition of Wasserstein Distance
We here provide the definition in the common case where probability density functions $p_{1}$ and $p_{2}$ are assumed to exist (formal definition using probability measures in Appendix A.3). Let $M\subseteq\mathbb{R}^{d}$, and $||\cdot||_{q}$ be the $q$-norm in $\mathbb{R}^{d}$. Then the Wasserstein-$q$ norm can be written as
$W_{q}(p_{1},p_{2})=\inf_{\gamma\sim\Gamma(p_{1},p_{2})}\left(\mathbb{E}_{x_{1},x_{2}\sim\gamma}\,||x_{1}-x_{2}||_{q}^{q}\right)^{\frac{1}{q}},$ (1)
where $\Gamma(p_{1},p_{2})$ is the set of all couplings, that is all possible “transportation plans”, between $p_{1}$ and $p_{2}$. $\gamma\in\Gamma(p_{1},p_{2})$ is a joint distribution over $(x_{1},x_{2})$ with respective marginals $p_{1}$ and $p_{2}$ over $x_{1}$ and $x_{2}$.
Sample-Based Wasserstein Distance
In practice, Eq. (1) is analytically solvable for only a few distributions. Therefore, Wasserstein distance is typically estimated from finite samples from $p_{1}$ and $p_{2}$. However, sample-based estimates of the Wasserstein distance are biased, and the convergence to the true Wasserstein distance is exponentially slower as the dimensionality of the distribution increases (Fournier & Guillin, 2015, Papp & Sherlock, 2022)
@@
Intuitively, if two given probability distributions are thought of as two piles of dirt, the Wasserstein distance measures the (minimal) cost of “transporting” one pile of dirt to another (Panaretos & Zemel, 2019). This analogy led to the Wasserstein distance also being known as the Earth Mover’s Distance. Suppose we have samples $\{x_{1},\ldots,x_{N}\}\subset\mathbb{R}^{d}$ sampled from a distribution $p_{1}$ and $\{y_{1},...,y_{N}\}\subset\mathbb{R}^{d}$ sampled from another distribution $p_{2}$. Given any distance metric between two vectors in $\mathbb{R}^{d}$, $D(\cdot,\cdot)$, we can construct the cost matrix $C$, as the matrix of pairwise distances between the samples $x_{i}$ and $y_{j}$:
@@
Recalling the Earth Mover’s Distance analogy, we want to map each $x_{i}$ to exactly one $y_{j}$, in such a way that the cost of doing so is minimized. The minimum transport map then defines the Wasserstein distance (for the metric $D$) between the two empirical distributions. Throughout this work, we employ the commonly used Euclidean metric, $L^{2}$, leading to the Wasserstein-2 and Sliced Wasserstein-2 distances. More precisely, we define a “transport map” to be a permutation matrix, $\pi\in\{0,1\}^{N\times N}$, which is a matrix with exactly one nonzero entry in each row. The entry $\pi_{ij}=1$ means that we transport the point $x_{i}$ to the point $y_{j}$. Then, finding the transport map that minimizes the overall cost can be stated as
@@
Figure 3: Computing Wasserstein distance. Two transport maps mapping the samples from a two-dimensional distribution $p_{1}$ (black) to samples from another distribution $p_{2}$ (blue), shown by arrows. The color of the arrow corresponds to the cost (Euclidean distance) between $x_{i}$ and $y_{i}$. (a) Randomly chosen transport map. (b) The optimal transport map, giving the smallest total cost. The total cost for the optimal map in (b) is the Wasserstein distance between these two sets of samples. Note that this schematic demonstrates transport maps for the non-sliced Wasserstein distance in two dimensions.
@@
Solving the optimal transport problem (Eq. (3)) with the Hungarian method has a time complexity of $O(N^{3})$ in the number of samples $N$ (although faster $O(N^{2}\log N$) approximations exist, see Peyré et al. 2017). However, in the special case where the data is one-dimensional, the Wasserstein distance can be calculated by sorting the two datasets, obtaining the order statistics $\{x_{(1)},..,x_{(N)}\}$ and $\{y_{(1)},...,y_{(N)}\}$ and computing the sum of the distances $\sum_{i}D(x_{(i)},y_{(i)})$. This has a time complexity of $O(N\log N)$. Thus, slicing the Wasserstein distance with one-dimensional projections becomes very efficient. While the value of the SW distance does not converge to the Wasserstein distance, even in the case of an infinite number of data samples and slices, the SW distance is a metric (in the mathematical sense) as long as $D$ is a metric on $\mathbb{R}^{d}$, and it acts as a lower bound to the Wasserstein distance (Nadjahi, 2021).
Limitations
When two high-dimensional distributions differ only along a small subset of directions, relying on one dimensional projections can become a significant limitation. This is because in high-dimensional spaces, most slices will be (almost) orthogonal to the few directions which distinguish the two distributions. In such a situation, the gain in efficiency from slicing can be impeded by the large number of slices we would need to reliably tell the two distributions apart. This limitation can be addressed by considering nonlinear (Kolouri et al., 2019) or other specific (Deshpande et al., 2019; 2018) slices. Furthermore, slicing may also be relaxed to other kinds of data-specific projections, such as Fourier features for stationary time series or locally-connected projections for images (Du et al., 2023, Cazelles et al., 2021). Finally, while sliced distances can overcome computational challenges of computing distances in higher dimensions, they typically inherit the limitations of the distance being sliced.